THE ANGLE OF ELEVATION OF A CLOUD FROM A POINT 60 M ABOVE A LAKE IS 30 DEGREE AND THE ANGLE OF DEPRESSION OF THE REFLECTION OF CLOUD IN THE LAKE IS 60 DEGREE. FIND THE HEIGHT OF THE CLOUD.

Grade:12th Pass

1 Answers

SAGAR SINGH - IIT DELHI
12 years ago

Dear student,

let height of cloud is H from ground then height of cloud from point is (H-60) ...

using tan30 = (H-60)/d      .........1                (d is the base of upper triangle formed)

 

image of cloud after reflection is formed at H depth below the water surface , so for lower triangle

 tan60 = (H + 60)/d          ...........2               (d is the base of lower triangle)

dividing 2 by 1

 H+60/H-60 = 3

 H + 60 = 3H - 180

 2H = 240

 H = 120m